# Boundary-value problems of functional differential equations with state-dependent delays

### v2

@@ Abstract
Keywords: functional differential equations, periodic solutions, boundary-value problems, collocation methods, state-dependent delay, numerical bifurcation analysis
@@ 1 Introduction
If the right-hand side of the differential equation depends on values of $y$ at other times than the current $t$, one speaks of functional differential equations, writing
$\dot{y}(t)$ $=G_{\mathrm{FDE}}(y_{t},p)\mbox{,}$ (1.1)
where the dependent variable is $y:\mathbb{R}\to\mathbb{R}^{n_{y}}$, $p\in\mathbb{R}^{n_{p}}$ are the problem parameters, and
$G_{\mathrm{FDE}}:C^{0}([-\tau_{\max},0];\mathbb{R}^{n_{y}})\times\mathbb{R}^{n_{p}}\to\mathbb{R}^{n_{y}}$
is a continuous nonlinear functional. We use $C^{\ell}(I;\mathbb{R}^{n})$ for the space of continuous (for $\ell=0$) or continuously differentiable (for $\ell>0$) functions from the interval $I\subset\mathbb{R}$ into $\mathbb{R}^{n}$. The subscript $t$ in $y_{t}$ denotes the time shift operator
$C^{0}(\mathbb{R};\mathbb{R}^{n_{y}})\times\mathbb{R}\ni(y,t)\mapsto y_{t}\in C([-\tau_{\max},0];\mathbb{R}^{n_{y}})\mbox{\ with\ }y_{t}(s)=y(t+s)\mbox{}$
(we may also simply write $\dot{y}(t)=G_{\mathrm{FDE}}(y(t+(\cdot)),p)$). Common challenging bifurcation analysis problems involving FDEs arise in optics due to transport delays (Seidel et al., 2022), in population dynamics due to non-zero maturation times (leading to implicitly defined threshold delays and integrals over the past (Gedeon et al., 2022; Diekmann et al., 2010), or in machining due to the effects of the machining tool on the surface from the previous revolution (Insperger et al., 2008).
@@
Complete tools for bifurcation analysis incorporating collocation for periodic orbits were developed and implemented by Engelborghs et al. (2002) (DDE-Biftool) and Szalai (2006) (knut), see also Roose and Szalai (2007) for a review. These tools permit an arbitrary number of discrete delays as part of $G_{\mathrm{FDE}}$, which may depend on the state and system parameters (for DDE-Biftool). A-posteriori convergence tests on examples suggested convergence orders equal to the degree $m$ of the polynomial pieces for the maximum norm of the error over interval $[0,1]$. One cannot expect better (e.g., superconvergence) as the interpolation uses the piecewise polynomial $y^{L}(\cdot)$ of degree $m$ when evaluating $G_{\mathrm{FDE}}$ at the collocation points (Engelborghs et al., 2001; Barton et al., 2006). Engelborghs and Doedel (2002) proved linear stability for collocation methods for time-periodic linear inhomogeneous FDEs with discrete constant delays. They pointed to “general stability theory for discretizations of nonlinear operator equations” for concluding (informally) convergence of the methods, referring to (Keller, 1975).
@@
$\dot{y}(t)=-y(t-p-y(t)).$ (1.3)
This DDE fits the general form (1.1) with $G_{\mathrm{FDE}}(y,p)=-y(-p-y(0))$, which is continuous on an open subset of $C^{0}([-\tau_{\max},0];\mathbb{R})\times\mathbb{R}$ for $\tau_{\max}>\frac{\pi}{2}$ (so, $n_{y}=n_{p}=1$). The Hopf bifurcation theorem ensures that this DDE has a family of small-amplitude periodic solutions of the form $y(t)=y_{0}\sin(\theta+t/T)+O(|y_{0}|^{2})$ with $0<|y_{0}|\ll 1$ and period $T=2\pi+O(|y_{0}|^{2})$ for $p=\pi/2+O(|y_{0}|^{2})$ and arbitrary $\theta$. See (Kuznetsov, 2004) for the classical Hopf bifurcation theorem, (Hale, 1977; Diekmann et al., 1995) for the version for FDEs with constant delay, and (Sieber, 2012) for its proof for FDEs with state-dependent delays. After rescaling the time interval to $[0,1]$ the unknown period $T$ appears explicitly as a parameter such that one is looking for $1$-periodic functions $y$, periods $T$, and parameters $p$ such that
@@
where we use $C^{\ell}_{\pi}$ for spaces of $\ell$ times continuously differentiable $1$-periodic functions. We observe that unknowns appear inside the arguments of $y$, which is itself unknown, such that application of the Fréchet derivative to ${G_{\mathrm{ts}}}$ reduces the regularity of the argument $y$ of ${G_{\mathrm{ts}}}$:
@@
$-\delta^{T}\left[y(t-p/T-y(t)/T)+y^{\prime}(t-p/T-y(t)/T)[p+y(t)]/T\right].$
Thus, the right-hand side ${G_{\mathrm{ts}}}$ is not differentiable or locally Lipschitz continuous if we consider ${G_{\mathrm{ts}}}$ as mapping $y\in C^{\ell}_{\pi}$ into $C^{\ell}_{\pi}$ for any $\ell\geq 0$. The review by Hartung et al. (2006) pointed out this lack of continuous differentiability and its consequences (see also (Cassidy et al., 2019)). For example, solutions for IVPs are not unique if one permits $C^{0}$ initial conditions. If the initial conditions are at least $C^{1}$ and compatible (Walther, 2003), then the dependence of IVP solutions on initial values is $C^{1}$ but results on higher regularity are missing. Large parts of the theory for FDEs as developed in textbooks by Hale (1977), Hale and Verduyn Lunel (1993) and Diekmann et al. (1995) relies strongly on continuous differentiability of ${G_{\mathrm{ts}}}$ acting on arguments $y\in C^{0}$, and are, thus, not applicable to problems such as (1.3).
@@
Similarly, discretization theory for boundary-value problems (BVPs) has been developed for nonlinearities ${G_{\mathrm{ts}}}$ that are Fréchet differentiable for $C^{0}$ arguments by Maset (2015b, a, 2016). Andò and Breda (2020) noted that for periodic BVPs with unknown period $T$ the rescaling by the unknown period $T$ introduces a state dependence of the deviating time arguments (namely on $T$, see the term $Ty(t-p/T\ldots)$ in (1.4)), even for FDE problems with constant delay. This causes a loss of differentiability of the problem’s nonlinearity with respect to the unknown $T$. Careful reanalysis of the general framework constructed by Maset (2016) showed that continuous differentiability with respect to the scalar variable $T$ is only needed in a single point, namely the assumed-to-exist solution of the BVP. Thus, Andò and Breda (2020) proved convergence of the routinely used methods in DDE-Biftool and knut for FDEs with constant delays, closing the gap left in the argument of Engelborghs and Doedel (2002).
The analysis of Andò and Breda (2020) leaves the question open how necessary special treatment of the unknown period $T$ is (a finite-dimensional part of the unknowns of the problem), or if convergence of numerical discretization for periodic BVPs can be proved without assuming Fréchet differentiability of the right-hand side. The review by Hartung et al. (2006) points to the appropriate generalized differentiability properties that are satisfied by the nonlinearities occuring in FDEs. As one can see in the right-hand side (1.5) of example (1.4) and its derivative (1.6), nonlinearities in FDEs are Fréchet differentiable $\ell$ times with respect to their arguments $y$ and $p$ only if $y\in C^{\ell}$. In (1.6) we also see that, while the derivative depends on $y^{\prime}$, it does not depend on $(\delta^{y})^{\prime}$. We use this property of restricted differentiability (we call it mild in Definition 2.1) to prove our central convergence result for discretizations of embedded periodic BVPs of FDEs.
@@ 2.1 Periodic BVPs for FDEs and mild differentiability
$y^{\prime}(t)$ $=TG_{\mathrm{FDE}}(y(t+(\cdot)/T),p),$ 00 0 $=R_{\mathrm{aff}}[y,T,p]\mbox{,}$ (2.1)
for continuous $1$-periodic $y:\mathbb{R}\to R^{n_{y}}$, period $T$, and parameter $p\in\mathbb{R}^{n_{p}}$, where $G_{\mathrm{FDE}}:C^{0}([-\tau_{\max},0];\mathbb{R}^{n_{y}})\times\mathbb{R}^{n_{p}}\to\mathbb{R}^{n_{y}}$ is continuous. The affine map $R_{\mathrm{aff}}:C^{0}_{\pi}\times\mathbb{R}^{n_{p}+1}\to\mathbb{R}^{n_{p}+1}$ defines constraints, making the system “square” when including the parameters $T$ and $p$ as unknowns. The constraints should also eliminate the shift symmetry that $y(\theta+(\cdot))$ is a solution of (1.3) for all $\theta$ whenever it is a solution for $\theta=0$. Nonlinearities such as $G_{\mathrm{FDE}}$ require the concept of mild differentiability to describe their regularity.
@@
    (restricted continuous differentiability) $G|_{C^{\ell}}$ is $\ell$ times differentiable for $\ell\leq{\ell_{\max}}$ (in particular, $G$ is continuous), and
@@
    (extendability) the map $C^{\ell}\times C^{\ell}\ni(u,\delta^{u})\mapsto D^{\ell}G(u)(\delta^{u})^{\ell}\in\mathbb{R}^{n_{G}}$ has a continuous extension to $C^{\ell}\times C^{\ell-1}$ for all $\ell\leq{\ell_{\max}}$.
@@
through the embedding of $\mathbb{R}^{n_{p}+1}$ into $C^{0}_{\pi}$ that treats a vector $(T,p)$ as the constant function $t\mapsto(T,p)$. The nonlinearity $(y,T,p)\mapsto-Ty(-p/T-y(0)/T)$ used in example (1.4) is an example of a mildly differentiable functional (to arbitrary degree).
@@
    (Mild differentiability) The right-hand side $G_{\mathrm{FDE}}$ in the DDE (2.1) is mildly differentiable to order ${\ell_{\max}}\geq 1$.
@@ 2.2 Convergence of solution of discretized BVP
For polynomial collocation the unknown function is a $1$-periodic continuous piecewise polynomial $y^{L}$ on a mesh given as $10=t_{0}<\ldots<t_{L}=$ 1. More precisely, $y^{L}$ is a polynomial of degree $m$ on $[t_{i-1},t_{i}]$ for all $i=1,\ldots,L$ in each of its $n_{y}$ components, $y^{L}$ is continuous, and $y(t)=y(t+1)$ for all $t$. Additional unknowns are the parameters $(T^{L},p^{L})$, resulting in $n_{y}mL+n_{p}+1$ unknowns overall. The system of algebraic equations,
00 0 $=(y^{L})^{\prime}(t_{i,j})-T^{L}G_{\mathrm{FDE}}(y^{L}(t_{i,j}+(\cdot)/T^{L}),p^{L}),$ 00 0 $=R_{\mathrm{aff}}[y^{L},T^{L},p^{L}]$ (2.2)
for $1\leq i\leq L$, $1\leq j\leq m$, evaluates the FDE at the collocation points $t_{i,j}=t_{i-1}+(t_{i}-t_{i-1})t_{\mathrm{c},j}$, where the points $t_{\mathrm{c},j}$ are the $m$ collocation points for degree $m-1$ on the interval $[0,1]$ for a sequence of orthogonal polynomials (e.g. of Gauss-Legendre or Chebyshev type). The strategy for adjusting approximation quality is a finite-element approach (Andò and Breda, 2020; Andò, 2020), keeping the degree $m$ bounded, and considering the limit $L\to\infty$, refining the mesh such that
$\max\{t_{i}-t_{i-1}\}\leq C_{\mathrm{msh}}/L$ (2.3)
for a bounded $C_{\mathrm{msh}}$ independent of $L$. In contrast, the strategy of keeping $L$ bounded and letting $m$ go to infinity is called the spectral element method in (Breda et al., 2005; Trefethen, 1996). Spectral methods promise exponential convergence, but require bounds on all derivatives of $y^{*}$. In the proofs of Lemmas 6.1 and 6.5 we rely on the boundedness of the degree $m$, but both Lemmas only require $G$ to be mildly differentiable once.
We can now state the following convergence theorem for the discretized BVP (2.2).
@@
Under Assumption 2.2 the discretized BVP (2.2) with (2.3) has a locally unique solution $x_{L}=(y^{L},T^{L},p^{L})$ near $x^{*}=(y^{*},T^{*},p^{*})$ for all sufficiently large $L$. The solution satisfies
$\|x_{L}-x^{*}\|_{0,1}:=\max\{\|y^{L}-y^{*}\|_{0,1},|T^{L}-T^{*}|,|p^{L}-p^{*}|\}=O\left(L^{-\min\{{\ell_{\max}},m\}}\right),$ (2.4)
@@
Andò (2020); Andò and Breda (2020) reformulate the BVP (2.1) as a fixed-point problem, following the general framework of Maset (2016). We modify this approach in section 3 to exploit the special structure present in periodic BVPs by constructing a fixed-point problem where the right-hand side is compact and maps spaces of periodic functions into spaces of periodic functions.
For this fixed-point problem the discretization corresponds to inserting a projection operator $\mathcal{P}_{L}$ onto the space of discontinuous piecewise polynomials of degree $m-1$ in between the nonlinearity and an integral operator. This formulation ensures that the space of numerical solutions $y^{L}$ consists of continuous $1$-periodic piecewise polynomials of degree $m$. Theorem 6.6 in section 6 establishes convergence for the discretized fixed-point problem.
We list basic properties of mildly differentiable functionals in Section 4. Validity of the chain rule ensures that $t\mapsto TG_{\mathrm{FDE}}(y(t+(\cdot)/T,p)$ is in $C^{\ell}$ if $y\in C^{\ell}$ and $G_{\mathrm{FDE}}$ is $\ell$ times mildly differentiable. Mild differentiability of $G_{\mathrm{FDE}}$ helps us in section 6 to establish consistence and stability of the numerical method and the smallness of the nonlinear terms, leading to the proof of Theorem 6.6 and, hence, Theorem 2.3.

### v3

@@ Abstract
Keywords: functional differential equations, delay differential equations, periodic solutions, boundary-value problems, collocation methods, state-dependent delay, numerical bifurcation analysis
@@ 1 Introduction
If the right-hand side of the differential equation depends on values of $y$ at times other than the current $t$, one speaks of functional differential equations, writing
$\dot{y}(t)$ $=G_{\mathrm{FDE}}(y_{t},p)\mbox{, where\quad}G_{\mathrm{FDE}}:C^{0}([-\tau_{\max},0];\mathbb{R}^{n_{y}})\times\mathbb{R}^{n_{p}}\to\mathbb{R}^{n_{y}}$ (1.1)
where the dependent variable is $y:\mathbb{R}\to\mathbb{R}^{n_{y}}$, $p\in\mathbb{R}^{n_{p}}$ are the problem parameters, and $G_{\mathrm{FDE}}$ is a continuous nonlinear functional. The most common example class for equations such as (1.1) are delay differential equations (DDEs). Our illustrative example (1.3) below is a DDE. We use $C^{\ell}(I;\mathbb{R}^{n})$ for the space of continuous (for $\ell=0$) or continuously differentiable (for $\ell>0$) functions from the interval $I\subset\mathbb{R}$ into $\mathbb{R}^{n}$. The subscript $t$ in $y_{t}$ denotes the time shift operator
[…]
(we may also write $\dot{y}(t)=G_{\mathrm{FDE}}(y(t+(\cdot)),p)$). Common challenging bifurcation analysis problems involving FDEs arise in optics due to transport delays (Seidel et al., 2022), in population dynamics due to non-zero maturation times (leading to implicitly defined threshold delays and integrals over the past (Gedeon et al., 2022; Diekmann et al., 2010)), or in machining due to the effects of the machining tool on the surface from the previous revolution (Insperger et al., 2008).
@@
Complete tools for bifurcation analysis incorporating collocation for periodic orbits were developed and implemented by Engelborghs et al. (2002) (DDE-Biftool) and Szalai (2006) (knut), see also (Roose and Szalai, 2007) for a review. These tools permit an arbitrary number of discrete delays as part of $G_{\mathrm{FDE}}$, which may depend on the state and system parameters (for DDE-Biftool). A-posteriori convergence tests on examples suggested convergence orders equal to the degree $m$ of the polynomial pieces for the maximum norm of the error over interval $[0,1]$. One cannot expect better (e.g., superconvergence) as the interpolation uses the piecewise polynomial $y^{L}(\cdot)$ of degree $m$ when evaluating $G_{\mathrm{FDE}}$ at the collocation points (Engelborghs et al., 2001; Barton et al., 2006). Engelborghs and Doedel (2002) proved linear stability for collocation methods for time-periodic linear inhomogeneous FDEs with discrete constant delays. They pointed to “general stability theory for discretizations of nonlinear operator equations” for concluding (informally) convergence of the methods, referring to (Keller, 1975). An alternative are methods based on series expansions of Chebyshev or Fourier type, so, fixing $L=1$, letting $m$ go to infinity, and choosing a projection of the right-hand side onto the truncated series up to $m$. The high-order a-posteriori convergence observed for these methods makes them excellent candidates for verified numerical computations. Gimeno et al. demonstrated how Chebyshev series truncations can be applied to some DDEs with a single state-dependent delay to prove existence of periodic orbits and their isochrons with verified numerics (Gimeno et al., 2023, 2021).
@@
$\dot{y}(t)=-y(t-p-y(t)),$ (1.3)
where $p\approx\pi/2$. This FDE fits the general form (1.1) with $G_{\mathrm{FDE}}(y,p)=-y(-p-y(0))$, which is continuous on an open subset of $C^{0}([-\tau_{\max},0];\mathbb{R})\times\mathbb{R}$ for $\tau_{\max}>\frac{\pi}{2}$ (so, $n_{y}=n_{p}=1$). The Hopf bifurcation theorem ensures that this FDE has a family of small-amplitude periodic solutions of the form $y(t)=y_{0}\sin(\theta+t/T)+O(|y_{0}|^{2})$ with $0<|y_{0}|\ll 1$ and period $T=2\pi+O(|y_{0}|^{2})$ for $p=\pi/2+O(|y_{0}|^{2})$ and arbitrary $\theta$. See (Kuznetsov, 2004) for the classical Hopf bifurcation theorem, (Hale, 1977; Diekmann et al., 1995) for the version for FDEs with constant delay, and (Sieber, 2012) for its proof for FDEs with state-dependent delays. After rescaling the time interval to $[0,1]$ the unknown period $T$ appears explicitly as a parameter we look for $1$-periodic functions $y$, periods $T$, and parameters $p$ such that
@@
where the subscript $(\cdot)_{\mathrm{ts}}$ stands for time shift and we use $C^{\ell}_{\pi}$ for spaces of $\ell$ times continuously differentiable $1$-periodic functions. We observe that unknowns appear inside the arguments of $y$, which is itself unknown, such that application of the derivative to $G_{\mathrm{ts}}$} reduces the regularity of the argument $y$ of $G_{\mathrm{ts}}$}:
@@
$-\delta^{T}\left[y(t-p/T-y(t)/T)+y^{\prime}(t-p/T-y(t)/T)[p+y(t)]/T\right]$
for a deviation $(\delta^{y},\delta^{T},\delta^{p})\in C^{\ell}_{\pi}\times\mathbb{R}\times\mathbb{R}$. Thus, the right-hand side $G_{\mathrm{ts}}$} is not differentiable or locally Lipschitz continuous if we consider $G_{\mathrm{ts}}$} as mapping $y\in C^{\ell}_{\pi}$ into $C^{\ell}_{\pi}$ for any $\ell\geq 0$. The review by Hartung et al. (2006) pointed out this lack of continuous differentiability and its consequences (see also Cassidy et al. (2019)). For example, solutions for IVPs are not unique if one permits $C^{0}$ initial conditions. If the initial conditions are at least $C^{1}$ and compatible (Walther, 2003), then the dependence of IVP solutions on initial values is $C^{1}$ but results on higher regularity are missing. Large parts of the theory for FDEs as developed in textbooks by Hale (1977), Hale and Verduyn Lunel (1993) and Diekmann et al. (1995) relies on continuous differentiability of $G_{\mathrm{ts}}$} acting on arguments $y\in C^{0}$, and are, thus, not applicable to problems such as (1.3).
@@
Similarly, discretization theory for boundary-value problems (BVPs) has been developed for nonlinearities $G_{\mathrm{ts}}$} that are differentiable for $C^{0}$ arguments by Maset (2015b, a, 2016). Indeed, it is tempting to assume differentiability of the nonlinearities and their discretized counterparts in FDE BVPs, because theorems on zeros of differentiable operators such as (Krasnosel’skii et al., 1972, Lemma 19.1) allow one to prove the uniqueness of the fixed point of the discretized problem straightforwardly. Andò and Breda (2020) pointed out that for periodic BVPs with unknown period $T$ the rescaling by the unknown period $T$ introduces a state dependence of the deviating time arguments (namely on $T$, see the term $Ty(t-p/T\ldots)$ in (1.4)), even for FDE problems with constant delay. This causes a loss of differentiability of the problem’s nonlinearity with respect to the unknown $T$. Careful reanalysis of the general framework constructed by Maset (2016) showed that continuous differentiability with respect to the scalar variable $T$ is only needed in a single point, namely the assumed-to-exist solution of the BVP. Thus, Andò and Breda (2020) proved convergence of the routinely used methods in DDE-Biftool and knut for FDEs with constant delays for the first time, closing the gap left in the argument of Engelborghs and Doedel (2002).
The analysis of Andò and Breda (2020) leaves the question open how necessary special treatment of the unknown period $T$ is (a finite-dimensional part of the unknowns of the problem), or if convergence of numerical discretization for periodic BVPs can be proved without assuming continuous differentiability of the right-hand side (also referred to as Fréchet differentiability in, e.g., (Ambrosetti and Prodi, 1995)) . The review by Hartung et al. (2006) points to the appropriate generalized differentiability properties that are satisfied by the nonlinearities occuring in FDEs. As one can see in the right-hand side (1.5) and its derivative (1.6), nonlinearities in FDEs are differentiable $\ell$ times with respect to their arguments $y$ and $p$ only if $y\in C^{\ell}$. In (1.6) we also see that, while the derivative depends on $y^{\prime}$, it does not depend on $(\delta^{y})^{\prime}$. We use this property of restricted differentiability (we call it mild in Definition 2.1) to prove our central convergence result for discretizations of embedded periodic BVPs of FDEs. In particular, our main result is the error estimate $O(L^{-\min\{\ell_{\max},m\}})$ for right-hand sides which are mildly differentiable to order $\ell_{\max}$, where $m$ is the fixed degree of the approximating piecewise polynomials and $L$ is the increasing size of the mesh.
@@ 2.1 Periodic BVPs for FDEs and mild differentiability
$y^{\prime}(t)$ $=TG_{\mathrm{FDE}}(y(t+(\cdot)/T),p),$ $0$ $=R_{\mathrm{aff}}[y,T,p]\mbox{,}$ (2.1)
for continuous $1$-periodic $y:\mathbb{R}\to\mathbb{R}^{n_{y}}$, period $T$, and parameter $p\in\mathbb{R}^{n_{p}}$, where $G_{\mathrm{FDE}}:C^{0}([-\tau_{\max},0];\mathbb{R}^{n_{y}})\times\mathbb{R}^{n_{p}}\to\mathbb{R}^{n_{y}}$ is continuous. The affine map $R_{\mathrm{aff}}:C^{0}_{\pi}\times\mathbb{R}^{n_{p}+1}\to\mathbb{R}^{n_{p}+1}$ defines constraints, making the system “square” when including the parameters $T$ and $p$ as unknowns. The constraints should also eliminate the shift symmetry that $y(\theta+(\cdot))$ is a solution of (1.3) for all $\theta$ whenever it is a solution for $\theta=0$. Nonlinearities such as $G_{\mathrm{FDE}}$ require the concept of mild differentiability to describe their regularity.
@@
    (restricted continuous differentiability) its restriction $G|_{C^{\ell}}:C^{\ell}\to\mathbb{R}^{n_{G}}$ is $\ell$ times differentiable for $\ell\leq{\ell_{\max}}$ (in particular, $G$ is continuous), and
@@
    (extendability) the map $C^{\ell}\times C^{\ell}\ni(u,\delta^{u})\mapsto D^{\ell}G(u)(\delta^{u})^{\ell}\in\mathbb{R}^{n_{G}}$ has a continuous extension to $C^{\ell}\times C^{\ell-1}$ for all $\ell\leq{\ell_{\max}}$, where $(\delta^{u})^{\ell}$ is the $\ell$-tuple $\delta^{u}\ldots\delta^{u}$ of arguments of the multilinear map $D^{\ell}G(u)$.
@@
through the embedding of $\mathbb{R}^{n_{p}+1}$ into $C^{0}_{\pi}$ that treats a vector $(T,p)$ as the constant function $t\mapsto(T,p)$. We list basic properties of mildly differentiable functionals in Section 4, which will be needed for our convergence analysis in Section 6. These include the validity of the chain rule, which ensures that $t\mapsto TG_{\mathrm{FDE}}(y(t+(\cdot)/T,p)$ is in $C^{\ell}$ if $y\in C^{\ell}$ and $G_{\mathrm{FDE}}$ is $\ell$ times mildly differentiable. The nonlinearity $(y,T,p)\mapsto-Ty(-p/T-y(0)/T)$ used in example (1.4) is an example of a mildly differentiable functional (to arbitrary degree). Definition 2.1 extends the classical “mild regularity assumption” made to obtain classical results for state-dependent FDEs (Walther, 2003; Hartung et al., 2006) to to arbitrary degree $\ell>1$. Alternative approaches taken in (Krisztin, 2003; Sieber, 2012) are equivalent but more involved to state and check in practical examples. All problems that can be formulated with standard numerical tools such as DDE-Biftool and ddesd available in MATLAB satisfy mild differentiability in the sense of Definition 2.1 if the problem coefficient functions are sufficiently regular. See (Humphries et al., 2015) for a demonstration of how DDE-Biftool can be used for FDEs with state-dependent delays and, hence, only mildly differentiable right-hand sides.
@@
    (Mild differentiability) The right-hand side $G_{\mathrm{FDE}}$ in the FDE (2.1) is mildly differentiable to order ${\ell_{\max}}\geq 1$.
@@ 2.2 Convergence of solution of discretized BVP
For polynomial collocation the unknown function is a $1$-periodic continuous piecewise polynomial $y^{L}$ on a mesh given as $0=t_{0}<\ldots<t_{L}=1$. More precisely, $y^{L}$ is a polynomial of degree $m$ on $t_{i-1},t_{i}]$ for all $i=1,\ldots,L$ in each of its $n_{y}$ components, $y^{L}$ is continuous, and $y(t)=y(t+1)$ for all $t$. Additional unknowns are the parameters $(T^{L},p^{L})$, resulting in $n_{y}mL+n_{p}+1$ unknowns overall. The system of algebraic equations,
$0$ $=(y^{L})^{\prime}(t_{i,j})-T^{L}G_{\mathrm{FDE}}(y^{L}(t_{i,j}+(\cdot)/T^{L}),p^{L}),$ $0$ $=R_{\mathrm{aff}}[y^{L},T^{L},p^{L}]$ (2.2)
for $1\leq i\leq L$, $1\leq j\leq m$, evaluates the FDE at the collocation points $t_{i,j}=t_{i-1}+(t_{i}-t_{i-1})t_{\mathrm{c},j}$, where the points $t_{\mathrm{c},j}$ are the $m$ collocation points for degree m1m-1 on the interval $[0,1]$ for a sequence of orthogonal polynomials (e.g. of Gauss-Legendre or Chebyshev type). The strategy for adjusting approximation quality is a finite-element approach (Andò, 2020; Andò and Breda, 2020), keeping the degree $m$ bounded, and considering the limit $L\to\infty$, refining the mesh such that $\max\{t_{i}-t_{i-1}\}\leq C_{\mathrm{msh}}/L$ for a bounded $C_{\mathrm{msh}}$ independent of $L$. In contrast, the strategy of keeping $L$ bounded and letting $m$ go to infinity is called the spectral element method in (Breda et al., 2005; Trefethen, 1996). Spectral methods promise exponential convergence, but require bounds on all derivatives of $y^{*}$. In the proofs of Lemmas 6.1 and 6.5 we rely on the boundedness of the degree $m$, but both Lemmas only require $G$ to be mildly differentiable once.
We can now state a convergence theorem for the discretized BVP (2.2).
@@
Under Assumption 2.2 the discretized BVP (2.2) with $\max\{t_{i}-t_{i-1}\}\leq C_{\mathrm{msh}}/L$ has a locally unique solution $x_{L}=(y^{L},T^{L},p^{L})$ near $x^{*}=(y^{*},T^{*},p^{*})$ for all sufficiently large $L$, satisfying
$\|x_{L}-x^{*}\|_{0,1}:=\max\{\|y^{L}-y^{*}\|_{0,1},|T^{L}-T^{*}|,|p^{L}-p^{*}|\}=O(L^{-\min\{{\ell_{\max}},m\}}).$ (2.3)
@@
Andò (2020); Andò and Breda (2020) reformulate the BVP (2.1) as a fixed-point problem, following the general framework in Maset (2016). We modify this approach in section 3 to exploit the special structure present in periodic BVPs by constructing a fixed-point problem where the right-hand side is compact and maps spaces of periodic functions into spaces of periodic functions.
For this fixed-point problem the discretization corresponds to inserting a projection operator $\mathcal{P}_{L}$ onto the space of discontinuous piecewise polynomials of degree m1m-1 in between the nonlinearity and an integral operator. This formulation ensures that the space of numerical solutions $y^{L}$ consists of continuous $1$-periodic piecewise polynomials of degree $m$. Theorem 6.6 in section 6 establishes convergence for the discretized fixed-point problem.
Mild differentiability of $G_{\mathrm{FDE}}$ helps us in section 6 to establish consistence and stability of the numerical method and the smallness of the nonlinear terms, leading to the proof of Theorem 6.6 and, hence, Theorem 2.3.